Jose, Paul, and Dan went fishing. it was a good day, and all three of them caught a fish. Dan's fish weighed four times as much as Paul's fish, and Jose's fish weighed half as much as Dan's fish. altogether, the three fish weighed 21 pounds. how much did each fish weigh?
Answer by mananth(16946)

Dan's fish weighed four times as much as Paul's fish,
and Jose's fish weighed half as much as Dan's fish.
the three fish weighed 21 pounds.
..
Let paul's fish weigh x lb
dan's fish = 4x
Jose's fish = 1/2 * 4x= 2x
..
x+4x+2x=21
7x=21
x=3 lbs . paul's fish
4x= 4*3=12 lbs Dan's fixh
2x = 2*3 = 6 lbs Jose's fish
..
CHECK
3+12+6 =21 lbs